if you lift a 50 lb barbell 6 foot off the floor, how much work do you do? by how much do you charge change the barbells potential energy? if you drop the barbell from 6-foot what is its kinetic energy just before it hits the floor?
Answer #1

m = 50 lbs = 22.679 kg

h = 6 ft = 1.8288 m

PE = Potential Energy

At Ground PE = 0

at 1.8288 m PE = mgh

PE = mgh

(a)

work done = PE = 22.679*1.8288*9.81

PE = 406.87 J

(b)

change the barbells potential energy

PE = 0 - 406.87

PE = 406.87 J

(c)

kinetic energy = potential energy = 406.87 J

velocity just before hitting ground

m*g*h = 1/2*m*v2

9.81*1.8288 J = 0.5.*v2

v = 6m/s